Project

General

Profile

Actions

RunPlan20161128 » History » Revision 35

« Previous | Revision 35/87 (diff) | Next »
George, Melissa, 04 December 2016 13:32


RunPlan20161128 to Dec 16th 2016

Scattering data with field, backup plan for straight tracks.

Daily schedule from 28/Nov to 16/Dec.

We expect to be running with 2V beamloss. Note Decay Solenoid is ON. ISIS is at 40+10 and 700 MeV

Note:

Date Nov 28 Nov. 29 Nov. 30 Dec. 1 Dec. 2 Dec. 3
MOM E. Overton M. Uchida M. Uchida M. Uchida M. Uchida M. Uchida
BLOC Y.Song Y.Song Y.Song Y.Song Y.Song R.Gamet
SOC D. Ragaram D. Ragaram D. Ragaram D. Ragaram D. Ragaram D. Ragaram
shifter 2 (08:00-16:00) Y.Song Y.Song Y.Song Y.Song Y.Song D.Bowring
Lead Shifter(08:00-16:00) P. Franchini P. Franchini P. Franchini P. Franchini P. Franchini P.Soler
Lead shifter (16:00-00:00) D. Ragaram D. Ragaram D. Ragaram D. Ragaram D. Ragaram R.Gamet
Shifter 2 (16:00-00:00) B.Freemire B.Freemire B.Freemire B.Freemire B.Freemire D.Maletic
Lead Shifter (00:00-08:00) G. Vankova G. Vankova G. Vankova G. Vankova G. Vankova B.Freemire 4-8
Shifter (00:00-08:00) J. Langlands J. Langlands J. Langlands J. Langlands J. Langlands P.Hanlet 12-4
Additional Personnel M.Courthold
Notes
Date Dec 4 Dec. 5 Dec. 6 Dec. 7 Dec. 8 Dec. 9
MOM M. Uchida M. Uchida M. Uchida M. Uchida M. Uchida M. Uchida
BLOC R.Gamet R.Gamet R.Gamet R.Gamet J.Greis J.Greis
SOC D. Ragaram D. Ragaram D. Ragaram D. Ragaram D. Ragaram D. Ragaram
Lead shifter (08:00-16:00) R.Gamet R.Gamet R.Gamet R.Gamet S.Wilbur S.Wilbur
Shifter 2 (08:00-16:00) D.Bowring D.Bowring D.Bowring D.Bowring J.Greis J.Greis
Lead shifter (16:00-00:00) P.Soler P.Soler P.Soler P.Soler D.Neuffer D.Neuffer
Shifter 2 (16:00-00:00) D.Maletic D.Maletic D.Maletic D.Maletic D.Maletic D.Maletic
Lead Shifter (00:00-08:00) M.Courthold M.Courthold P.Kyberd P.Kyberd A.Bross A.Bross
Shifter (00:00-08:00) B.Freemrire B.Freemrire P.Hanlet P.Hanlet B.Freemrire B.Freemrire
Additional Personnel
Notes

Plan for 2/Dec-5/Dec:

Dec 3-4th: From midnight: Finish H4a current run. Then cycle through 2 hours each of H6a-j and repeat.
Dec 5th: 10am Cooling channel magnet setting changes. Then Begin 2 hrs of each H7 and H8 until complete at new settings.

Data to take

Field On Data

Do 2 hours of each H4 task.

Task H4 - Emittance Reduction!!

Task Momentum # Triggers Beamline Tag CoolingChannel Tag Similar runs
H4a 140 1e6 TOF1 3-140+M3-Test2 2016-04-01.2
H4b 140 1e6 TOF1 6-140+M3-Test2 2016-04-01.2
H4c 140 2e6 TOF1 10-140+M3-Test3 2016-04-01.2

Task H6 - Diffuser Scan

The beamline and cooling channel settings remain the same and the diffuser mus be changed manually from the diffuser GUI for each new setting! Be sure to do this AFTER entering the beamline optics, or it will reset the diffuser. (Please note there is no run H6d.

Task Momentum # Triggers Beamline Tag CoolingChannel Tag Diffuser Setting Closed Iris Similar runs
H6a 170 5e5 TOF1 3-170+M3-Test1 2016-04-01.2 1 1 8498
H6b 170 5e5 TOF1 3-170+M3-Test1 2016-04-01.2 2 2 8497
H6c 170 5e5 TOF1 3-170+M3-Test1 2016-04-01.2 4 3 8495-6
H6e 170 5e5 TOF1 3-170+M3-Test1 2016-04-01.2 6 2+3 8499
H6f 170 5e5 TOF1 3-170+M3-Test1 2016-04-01.2 7 1+2+3 8502
H6g 170 5e5 TOF1 3-170+M3-Test1 2016-04-01.2 8 4 8494
H6h 170 5e5 TOF1 3-170+M3-Test1 2016-04-01.2 9 1+4 8500
H6i 170 5e5 TOF1 3-170+M3-Test1 2016-04-01.2 12 3+4 8501
H6j 170 5e5 TOF1 3-170+M3-Test1 2016-04-01.2 15 1+2+3+4 8503

Task H7 - Emittance Reduction!!

Task Momentum # Triggers Beamline Tag CoolingChannel Tag Similar runs
H7a 140 1e6 TOF1 3-140+M3-Test2 2016-04-1.5 H4a
H7b 140 1e6 TOF1 6-140+M3-Test2 2016-04-1.5 H4b
H7c 140 2e6 TOF1 10-140+M3-Test3 2016-04-1.5 H4c

Task H8 - Magnet alignment

Task Momentum # Triggers Beamline Tag CoolingChannel Tag Similar runs
H8a 170 5e5 TOF1 3-170+M3-Test1 2016-04-1.5 H5a
H8b 200 5e5 TOF1 3-200+M3_Test1 2016-04-1.5 H5b
H8c 240 5e5 TOF1 3-240+M3_Test1 2016-04-1.5 H5c

Straight track data (fallback)

Task S2: LiH scattering data.

Task Momentum # Triggers Beamline Tag CoolingChannel Tag Similar runs
S2a 140 1M @ TOF2 3-140+M3-Test2
S2b 200 ? 6-200+M0

Task S2a - 10M triggers at TOF1, 33 triggers/spill = 90 Hours data taking (15 shifts). Lots, but will make a
start anyway in case there is opportunity in later data taking for more running.

Physics post-processing

Step4_Pionic_beam_Tags

h2: Finished data sets:

h4. Task H5 - Magnet alignment (DONE)

Task Momentum # Triggers Beamline Tag CoolingChannel Tag Similar runs Final Event tally
H5a 170 5e5 TOF1 3-170+M3-Test1 2016-04-01.2 9.55e6
H5b 200 5e5 TOF1 3-200+M3_Test1 2016-04-01.2 8.92e6
H5c 240 5e5 TOF1 3-240+M3_Test1 2016-04-01.2 7.93e6

Updated by George, Melissa almost 7 years ago · 35 revisions